# Searching for Periodicity in FRB 20240114A

### v1

@@ Abstract
FRB 20240114A is extraordinarily active, and therefore presents an opportunity to search for the periodicity predicted by magnetar models of Fast Radio Bursts (FRB). Zhang et al. (2025) observed 11,553 bursts, including 3196 on MJD 60381 (March 12, 2024). We find no significant peak in the periodogram of those bursts, which occur within 15628 s, short enough that even with a characteristic slowing age of 1 year a signal with period $\geq 0.1\,$s it would not significantly dephase within the observation. Introducing modulation artificially shows that an amplitude of 0.15 would have been detected robustly. Implications for magnetar models of FRB are briefly considered.
@@
I Introduction
The most popular model of Fast Radio Bursts (FRB) (Zhang, 2023) attributes them to magnetars, hypermagnetized neutron stars. Magnetars were first proposed (Katz, 1982; Duncan & Thompson, 1992) as models of Soft Gamma Repeaters (SGR; initially considered to be gamma-ray bursts), a model that was verified when rapid spindown between outbursts confirmed their extraordinary magnetic fields. This model makes one obvious prediction: that FRB, or the level of FRB activity are periodic with the neutron star’s rotational period. This prediction has never been verified for any FRB, despite evidence Niu, Li, Yang et al. (2025); Wang, Lan, Zhao et al. (2025) that at least one FRB is inside an expanding supernova remnant, as would be exptected in a magnetar model.
@@ II.1 RRAT periodicity
II.2 Periodic Modulation of Activity
@@
III Parameters and Phase Drift
@@ IV Results
Results for frequencies (in cycles/s, not radians) up to 10 Hz are shown in Fig. 1. Extending the search to frequencies up to 100 Hz found no $P(\omega)$ larger than the largest found below 10 Hz (at a frequency of 2.19 Hz). The slight rise of the periodogram to the lowest frequencies indicates a modest variation of the rate of bursts through the 156280 s (4.34 hours) of observation.
Figure 1: Periodogram of 3196 bursts of FRB 20240114A in a 15628 s observation March 12, 2024 (MJD 60381), data from Zhang et al. (2025). Extension of the calculation to frequencies of 100 Hz (implying a neutron star with $\mu_{33}\ll 1$) produced no higher peaks than shown here.
V Frequency Derivative
@@
VI Sensitivity
In order to determine the sensitivity of the periodogram to periodic variation of the rate of FRB we introduce such variation artificially. Bursts are removed from the data with a probability $M\cos{(\omega_{m}t)}$, where $M$ is the modulation amplitude and $\omega_{m}$ the (angular) modulation frequency. We choose a frequency (in cycles) of 1 Hz; the weak dependence of the periodogram on frequency shows that the sensitivity depends only weakly on frequency.
When there are a large number of frequencies considered the clearest way of searching for periodicity is with a bar graph of the distribution of amplitudes (Nowak & Katz (2026) did this in a search for periodicity in the X-ray emission of Sco X-1 in which there were $2^{29}\approx 5.37\times 10^{8}$ frequencies and amplitudes, that could not be readily and legibly plotted). The results are shown in Fig. 2 for several values of $M$.
@@ VI Sensitivity
VII Discussion
Eq. 3 shows that large negative $\dot{\omega}$ only occurs for neutron stars that are both young and comparatively weakly magnetized; in other words, not magnetars. Quantitatively, our failure to find periodicity (for any possible $\omega$) with ${\dot{\omega}}\in[10^{-6},10^{-6}]$ s-2 indicates $\mu_{33}\leq 0.57A_{y}^{-3/2}$. If A>1A>1\,y, the implied $\mu_{33}$ would be below the range considered to characterize magnetars. Larger $|\dot{\omega}|$ outside the range calculated would require $\omega$ so large as to be inconsistent with Eq. 1 and the assumptions of magnetar fields and A>1A>1\,y.
@@ VII Discussion
VIII Conclusion
The magnetar model of FRB makes no quantitative prediction of the magnitude of modulation of the event rate at the magnetar’s rotation frequency. We therefore cannot say that our upper bound of 15% modulation in FRB 20240114A disproves this model. The bound is much lower than the modulation of radiated intensities of known magnetic neutron stars (radio pulsars, accreting neutron star X-ray sources, SGR and X-ray emitting magnetars), that are generally ${\cal O}(100\%)$, but without a quantitative theory of magnetar FRB any inference must be uncertain.
Magnetar models are consistent with tight upper limits on rotational modulation of the rate of detected FRB if either
@@ VIII Conclusion
Acknowldgements
I thank T. Piran for discussions and the Hebrew University for hospitality during the completion of this work.

### v2

@@ Abstract
FRB 20240114A is extraordinarily active, and therefore presents an opportunity to search for the periodicity predicted by magnetar models of Fast Radio Bursts (FRB). [Zhang et al. (2025)] observed 11,553 bursts, including 3196 on MJD 60381 (March 12, 2024). We find no significant peak in the periodogram of those bursts, which occur within 15628 s, short enough that even with a characteristic slowing age of 1 year a signal with period $\geq 0.1\,$s it would not significantly dephase within the observation. Introducing modulation artificially shows that an amplitude of 0.15 would have been detected robustly.
keywords:
@@
1 Introduction
The most popular model of Fast Radio Bursts (FRB) attributes them to magnetars, hypermagnetized neutron stars. Magnetars were first proposed (Katz, 1982; Duncan & Thompson, 1992) as models of Soft Gamma Repeaters (SGR; initially considered to be gamma-ray bursts), a model that was verified when rapid spindown between outbursts confirmed their extraordinary magnetic fields. This model makes one obvious prediction: that FRB, or the level of FRB activity are periodic with the neutron star’s rotational period. This prediction has never been verified for any FRB, despite evidence Niu, Li, Yang et al. (2025); Wang, Lan, Zhao et al. (2025) that at least one FRB is inside an expanding supernova remnant, as would be expected in a magnetar model.
@@ 2.1 RRAT periodicity
2.2 Periodic Modulation of Activity
@@
3 Parameters and Phase Drift
@@ 4 Results
Results for frequencies (in cycles/s, not radians) up to 10 Hz are shown in Fig. 1. Extending the search to frequencies up to 100 Hz found no $P(\omega)$ larger than the largest found below 10 Hz (at a frequency of 2.19 Hz). The slight rise of the periodogram at the lowest frequencies indicates a modest slow variation of the rate of detections through the 15628 s (4.34 hours) of observation. This may be intrinsic, or it may be the consequence of a variation of instrumental sensitivity with zenith angle.
Figure 1: Periodogram of 3196 bursts of FRB 20240114A in a 15628 s observation March 12, 2024 (MJD 60381), data from Zhang et al. (2025). Extension of the calculation to frequencies of 100 Hz (implying a neutron star with $\mu_{33}\ll 1$ unless its age is $\ll 1\,$y) produced no higher peaks than shown here.
5 Frequency Derivative
@@
6 Sensitivity
In order to determine the sensitivity of the periodogram to periodic variation of the rate of FRB we introduce such variation artificially. Bursts are removed from the data with a probability $M\cos{(\omega_{m}t)}$, where $M$ is the modulation amplitude and $\omega_{m}$ the (angular) modulation frequency. We choose a frequency (in cycles) of 1 Hz; the weak dependence of the periodogram (Fig. 1) on frequency shows that the sensitivity depends only weakly on frequency.
When there are a large number of frequencies considered the clearest way of searching for periodicity is with a bar graph of the distribution of amplitudes (Nowak & Katz (2026) did this in a search for periodicity in the X-ray emission of Sco X-1 in which there were $2^{29}\approx 5.37\times 10^{8}$ frequencies and amplitudes, that could not be readily and legibly plotted). Our results for FRB 20240114A are shown in Fig. 2 for $M=0$ (the actual data) and for several larger values of imposed modulation $M$ at 1 Hz, a plausible spin frequency of a young magnetar. Fig. 1 shows that the noise level, and therefore sensitivity to periodic variation, are nearly independent of the period of variation.
@@ 6 Sensitivity
7 Periodicity?
In the model of steady dipole spindown, Eq. 3 shows that large negative $\dot{\omega}$ only occurs for neutron stars that are both young and comparatively weakly magnetized; in other words, not magnetars. Quantitatively, our failure to find periodicity (for any possible $\omega$) with ${\dot{\omega}}\in[10^{-6},10^{-6}]$ s-2 indicates $\mu_{33}\leq 0.57A_{y}^{-3/2}$. If A>1A>1\,y, the implied $\mu_{33}$ would be below the range considered to characterize magnetars. Larger $|\dot{\omega}|$ outside the range calculated would require $\omega$ so large as to be inconsistent with Eq. 1 and the assumptions of magnetar fields and A>1A>1\,y.
@@ 7 Periodicity?
8 Conclusion
The magnetar model of FRB makes no quantitative prediction of the magnitude of modulation of the event rate at the magnetar’s rotation frequency. We therefore cannot say that our upper bound of 15% modulation in FRB 20240114A disproves this model. The bound is much lower than the modulation of radiated intensities of known magnetic neutron stars (radio pulsars, accreting neutron star X-ray sources, magnetars in their quiescent X-ray emitting state (Anomalous X-Ray Pulsars) and in their Soft Gamma Repeater outbursts), that are generally ${\cal O}(50\%)$, but without a quantitative theory of magnetar FRB any inference must be uncertain.
If the magnetar is not triaxial, magnetar models may still be consistent with tight upper limits on rotational modulation of the rate of detected FRB if either
@@ 8 Conclusion
The suggestion that repeating FRB have rotational and magnetic axes aligned with the line of sight, while apparent non-repeaters do not, implies thae we only observe a tiny fraction of the bursts of apparent non-repeaters but that we also observe a similarly tiny fraction of the repeaters.
@@
Data Availability
This theoretical study produced no new data.
Acknowledgments
I thank T. Piran for discussions and the Hebrew University, Jerusalem, Israel for hospitality during the completion of this work.